MW SPECTROSCOPY COUPLED WITH ULTRAFAST UV LASER VAPORIZATION: RIBOSE FOUND IN THE GAS PHASE

EMILIO J. COCINERO, PATRICIA ECIJA, FRANCISCO J. BASTERRETXEA, JOSÉ A. FERNÁNDEZ, FERNANDO CASTANO, Departamento de Química Física, Facultad de Ciencia y Tecnología, Universidad del País Vasco (UPV-EHU), Ap. 644, E-48080 Bilbao, Spain; ALBERTO LESARRI, Departamento de Química-Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47011 Spain; JENS-UWE GRABOW, Institut für Physikalische Chemie, Lehrgebiet A, Universität Hannover, Callinstraße. 3A, D-30167 Hannover, Germany.

Sugars are aldoses or ketoses with multiple hydroxy groups which have been elusive to spectroscopic studies. Here we report a rotational study of the aldopentose ribose. According to any standard textbook aldopentoses can exhibit either linear forms, cyclic five-membered (furanose) structures or six-membered (pyranose) rings, occurring either as α- or βanomers depending on the orientation of the hydroxy group at C-1 (anomeric carbon). β-Furanose is predominant in ribonucleosides, RNA, ATP and other biochemically relevant derivatives, but is β-furanose the native form also of free ribose? Recent condensed-phase X-ray^{a} and older NMR^{b} studies delivered conflicting results. In order to solve this question we conducted a microwave study on D-ribose that, owing to ultrafast UV laser vaporization^{c}, has become the first C-5 sugar observed with rotational resolution. The spectrum revealed six conformations of free ribose, preferentially adopting βpyranose chairs as well as higher-energy α-pyranose forms. The method also allowed for unambiguous distinction between different orientations of the hydroxy groups, which stabilize the structures by cooperative hydrogen-bond networks. No evidence was observed of the α-/β-furanoses or linear forms found in the biochemical derivatives.

a i) D. Šišak, L. B. McCusker, G. Zandomeneghi, B. H. Meier, D. Bläser, R. Boese, W. B. Schweizer, R. Gylmour and J. D.

Dunitz Angew. Chem. Int. Ed. 49, 4503, 2010. ii) W. Saenger Angew. Chem. Int. Ed. 49, 6487, 2010.

b i) M. Rudrum, and D. F. Shaw, J. Chem. Soc. 52, 1965. ii) R. U. Lemieux and J. D. Stevens Can. J. Chem. 44, 249, 1966.

iii) E. Breitmaier and U. Hollstein Org. Magn. Reson. 8, 573, 1976.

c E. J. Cocinero, A. Lesarri, P. Écija, F. J. Basterretxea, J. U. Grabow, J. A. Fernández and F. Castaño Angew. Chem. Int. Ed.

in press: DOI: 10.1002/anie.201107973, 2012.